# Advances in Tabulating Carmichael Numbers

### v2

@@ Abstract.
We report that there are $49679873$ Carmichael numbers less than $10^{22}$ which is an order of magnitude improvement on the previous bound. The tabulation finds Carmichael numbers of the form $n=Pqr$ and is bifurcated by the size of $P$ with respect to the tabulation bound $B$. For $P<7\cdot 10^{7}$, we found $35985331$ Carmichael numbers and $1202914$ of them were less than $10^{22}$. When $P>7\cdot 10^{7}$, we found $48476959$ Carmichael numbers less than $10^{22}$. We provide a comprehensive overview of both cases of the algorithm. For the large case, we show and implement asymptotically faster ways to tabulate compared to the prior tabulation. We also provide an asymptotic estimate of the cost of this algorithm. It is interesting that Carmichael numbers are worst case inputs to this algorithm. So, providing a more robust asymptotic analysis of the cost of the algorithm would likely require resolution of long-standing open questions regarding the asymptotic density of Carmichael numbers.
@@ 7. Asymptotic counts of Preproducts
We believe, as a guide, that the cost of completing a given preproduct $Pq$ is a polynomial-time algorithm, and that the bound in Theorem 8 serves as a guide for the count of preproducts constructed. A better guide might tell us the distribution of preproducts in $[0,B]$. In Section 5, we saw that the larger a preproduct was, the more likely it was to be an “easy case.” So, both the number of difficult problems and their distribution is important to establish sharper asymptotic bounds.
@@ 8. Implementation, statistics, and conclusion
$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr d&3&4&5&6&7\\ \hline\cr C_{d}(10^{22})&420658&2327421&1370257&6034046&14056372\\ \hline\cr\hline\cr 8&9&10&11&12&13\\ \hline\cr 16005645&8939434&2347828&262818&10018&55\\ \hline\cr\end{array}$
Each of these numbers was proven to be a Carmichael number with a separate check program, and we confirmed that the numbers up to $10^{21}$ matched Pinch’s most recent tabulation.
@@ 8.2. Large Preproduct Information
$\begin{array}[]{|c|c|c|c|c|}\hline\cr 4&5&6&7&8\\ \hline\cr 100164&1024718&5769410&14017455&16005059\\ \hline\cr\hline\cr 9&10&11&12&13\\ \hline\cr 8939434&2347828&262818&10018&55\\ \hline\cr\end{array}$

### v3

@@ Abstract.
We report that there are $49679870$ Carmichael numbers less than $10^{22}$ which is an order of magnitude improvement on Richard Pinch’s prior work. We find Carmichael numbers of the form $n=Pqr$ using an algorithm bifurcated by the size of $P$ with respect to the tabulation bound $B$. For $P<7\cdot 10^{7}$, we found $35985331$ Carmichael numbers and $1202914$ of them were less than $10^{22}$. When $P>7\cdot 10^{7}$, we found $48476956$ Carmichael numbers less than $10^{22}$. We provide a comprehensive overview of both cases of the algorithm. For the large case, we show and implement asymptotically faster ways to tabulate compared to the prior tabulation. We also provide an asymptotic estimate of the cost of this algorithm. It is interesting that Carmichael numbers are worst case inputs to this algorithm. So, providing a more robust asymptotic analysis of the cost of the algorithm would likely require resolution of long-standing open questions regarding the asymptotic density of Carmichael numbers.
@@ 7. Asymptotic counts of Preproducts
We believe, as a guide, that the cost of completing a given preproduct $Pq$ is a polynomial-time algorithm, and that the bound in Theorem 8 serves as a guide for the count of preproducts constructed. A better guide might tell us the distribution of preproducts in $[0,B]$. In Section 5, we saw that the larger a preproduct was, the more likely it was to be an easy case. So, both the number of difficult problems and their distribution is important to establish sharper asymptotic bounds.
@@ 8. Implementation, statistics, and conclusion
$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr d&3&4&5&6&7\\ \hline\cr C_{d}(10^{22})&420658&232742&1370257&6034046&14056367\\ \hline\cr\hline\cr 8&9&10&11&12&13\\ \hline\cr 16005646&8939435&2347828&262818&10018&55\\ \hline\cr\end{array}$
Each of these numbers was proven to be a Carmichael number with a separate check program, and we confirmed that the numbers up to $10^{21}$ matched Pinch’s most recent tabulation. Testing was also aided by a comparison against an unpublished tabulation performed by Claude Goutier.
@@ 8.2. Large Preproduct Information
$\begin{array}[]{|c|c|c|c|c|}\hline\cr 4&5&6&7&8\\ \hline\cr 100164&1024718&5769410&14017450&16005060\\ \hline\cr\hline\cr 9&10&11&12&13\\ \hline\cr 8939435&2347828&262818&10018&55\\ \hline\cr\end{array}$